40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 837     Entry time: Thu Aug 14 19:35:54 2008
Author: Jenne 
Type: Summary 
Category: IOO 
Subject: PRM in the chamber, ready to pump! 
Rob, Yoichi, Jenne, Steve

Summary: Everything is back in the chamber, we just need to put on the big doors, and start pumping in the morning.

After letting the PRM's standoff epoxy cure overnight, it was time to put the optic back in the BS Chamber. Rob put the optic cage back in the chamber, as close to the guide points that Rana had placed as possible. A handy technique was discovered for pushing the cage into place: put a long screw into the table, leaving an inch or so above the table, then use that as a push-off point so that you can push the base of the cage with your thumb. According to Rob, this is probably just about as effective as using a pusher-screw.

The guides were helpful in getting the PRM back to its original position, but one of them was placed in such a way that it could move when pushed against. The clamp that was used as a guide point was placed with one of the screws half on the edge of a hole, so that when the cage was pushed against the guide point, that screw could wiggle around, causing the clamp to rotate thus no longer being a definite guide point.

Just after putting the PRM in place, Rob found the standoff that had gone missing. (see elog #835)

Once the PRM was back in place, we put the OSEMs back, and reinstalled the satellite boxes that had been removed (PRM's, which Ben has fixed - an op-amp was blown, and BS's, which we used over in the clean room with the spare OSEMs). We found a problem with the LR PRM OSEM reading on Dataviewer. It was saturating when the OSEM was just sitting on the table, with nothing between the LED and the sensor. We measured the output from the satellite box with the octopus cables, and measured 2.3 volts, which is too much for the DAQ. It seems fine when we install it in the cage, and the magnet is blocking part of the light. We should investigate the gain of the satellite box when convenient. This is not something that needs to be done prior to pump-down. Also, when we put an allen wrench to block the light while checking which OSEM was which, we noticed that the Dataviewer reading would go down to -2V, then come back to 0V when the light was completely blocked. This may be some incorrect compensation for some whitening. Again, we should look into this, but it is not terribly time-sensitive.

Once the OSEMs were centered, we tried to turn on the damping for the PRM. This was successful, so we are confident that we have put all of the OSEMs back in their correct places.

We found that we were easily able to get the PRM's oplev back on the QPD, so we ~centered the oplev, and then centered all of the PRM's OSEMs. This assumes that the oplev was in a good place, but I think we've determined that this is the case.

We did the same thing for the SRM and the BS, to check the OSEM values before we close up for good. We found that some of the SRM OSEMs were reading low (magnet too far in), and that all of the BS OSEMs were low, perhaps as if the table were tilted a tiny bit after removing and replacing the weight of the PRM. We recentered all of the OSEMs for both of these optics.

We checked that all of the pigtails for the PRM OSEMs were anchored to the PRM cage using some copper wire as tie-downs.

We checked that all of the earthquake stops were within 1mm or so of each of the 3 optics in the BS chamber. The SRM's earthquake stops were fairly far out. One of the bottom ones was far enough that when Yoichi turned it the wrong way (accidentally), it fell out. He put it back in, and adjusted all of the earthquake stops appropriately. This 1mm distance comes from Seji, and the specs for the optics' cages.

We did a look-through of the chamber, and took out all of the tools, and other things that were not bolted down to the table.

We have left the damping of the PRM off for the night.

To do: put the doors back on, and start the pump down.
ELOG V3.1.3-